summary for week of 7/2-7/8/99

Ed Blucher (blucher@hep.uchicago.edu)
Sat, 10 Jul 1999 11:07:09 -0500

Statistics for Week of July 2-8, 1999

Runs: 13791-13838

Vac 2pi0 this week (Online w/cuts) 286K

Vac 2pi0 for 1999 run 910K

Hours of beam 101.5

Hours of datataking 92.4

Detector efficiency 0.91

Comments:

Problems during the last week include
- ET power supply failure
- UCLA Shaper/Splitter power supply failure
- ktev1 "fire"
- 1 tape drive failure (drv 3 on plane 2)

The SEM calibration was changed during run 13838. New SEM readings are
reduced by 21% compared to previous readings.

Beginning with run 13829, the drift chamber TDC window was shifted
by 96 ns to allow us to see more early hits. With this shift, accidental
events should include all hits needed to simulate inefficiencies resulting
from early accidental activity.

Aside from a single day at higher intensity, all data collected so far has been
at an intensity of ~4.1 E12 (using the new SEM calibration). At the Friday
meeting, we decided to take a week of data at higher intensity (~7E12).
There will be more discussion of the intensity plan for the rest of the run
at next Friday's meeting.